# Active Learning with Variational Quantum Circuits for Quantum Process Tomography

### v1

@@ Introduction
Numerical Experiments
In this section, we validate our algorithms through numerical simulations on the quantum processes of randomly generated circuits. Specifically, we conduct simulations with up to five qubits to evaluate the scalability of our algorithms.
Experimental Settings
Unknown Unitary Process.
We consider the quantum process of randomly generated circuits, as shown in Fig. 3. The circuit begins and ends with Hadamard gates applied to each qubit. Each two-qubit layer is counted as a depth, including controlled-phase (CZ) gates alternating between odd and even depths, and randomly chosen single-qubit gate ($T$, $\sqrt{X}$, or $\sqrt{Y}$). It is noted that such randomly generated quantum circuits are hard for efficient simulation on a classical computer (Boixo et al. 2018; Bouland et al. 2019).
Figure 3: Randomly generated quantum circuit
Baseline and Performance Metrics.
We consider a baseline random (RAND) that selects a subset of quantum states randomly. Additionally, we use similarity between the target process $\mathcal{U}$ and the approximate reconstruction $\mathcal{C}(\boldsymbol{\theta})$ to measure performance, defined as
$\textnormal{similarity}\left(\mathcal{U},\mathcal{C}(\boldsymbol{\theta})\right)=1-\frac{\|\mathcal{C}(\boldsymbol{\theta})-\mathcal{U}\|_{F}}{2\|\mathcal{U}\|_{F}}.$
Here, $\|\cdot\|_{F}$ denotes the Frobenius norm of a matrix. The $\text{similarity}(\mathcal{U},\mathcal{C}(\boldsymbol{\theta}))=1$ means $\|\mathcal{C}(\theta)-\mathcal{U}\|_{F}=0$, in which case $\mathcal{U}$ can be perfectly reconstructed from $\mathcal{C}(\boldsymbol{\theta})$.
@@
Other Experimental Parameters.
We conduct simulations with different numbers of qubits to evaluate the scalability of our algorithms. For an $n$-qubit quantum process ($n$ ranges from 2 to 5), we set the VQC depth to 3, 5, 7 and 8 to learn the corresponding quantum processes. The number of committee members in QBC and the number of ensemble models in EMCM is empirically set to be 6. To avoid random fluctuation, each experiment is repeated 100 times and the averaged metrics are reported.
Experimental Results
Figure 4: The comparison results with the metric of average similarity.
Figure 5: The comparison results with the metric of average improvement.
Similarity Metric.
The results related to similarity are shown in Fig. 4. We observe the following.
@@
  • QBC and EMCM significantly outperformed RAND across all numbers of labeled quantum states. GS initially performs worse than RAND when only a few quantum states are labeled. However, as the number of labeled quantum states increases, GS eventually outperforms RAND. A possible explanation for the phenomena is as follows. At the initial steps, GS might select some outliers to maximize the distance from the previously labeled quantum states, which leads to a decrease in performance. Since the number of outliers is usually very small, their impact diminishes as more quantum states are labeled. It is worth noting that as the system size increases, reconstructing the quantum process requires a significantly larger number of labeled quantum states, which helps mitigate the impact of outliers. Therefore, GS remains a promising approach, especially in large-scale quantum systems.
Improvement Metric.
The results related to improvement are shown in Fig. 5. We observe the following.
  • When $n$ ranges from 2 to 4, the improvement of QBC and EMCM initially increases and then decreases. This is because, with sufficient labeled quantum states, even randomly selected quantum states can lead to well-trained models. Meanwhile, the improvement of GS initially decreases, which is consistent with our previous analysis, as GS may select outliers in the initial steps.
  • When $n$ increases to 5, the improvement of all our algorithms does not exhibit a significant decreasing trend. This is because, as the system size increases, RAND requires more labeled quantum states to train a high quality model. Despite a large number of quantum states already being labeled, which supports good performance for our algorithms, this is still insufficient for RAND, resulting in consistently poor performance. Thus, our algorithms are crucial in large-scale quantum systems, where RAND always fails to obtain enough labeled quantum states to train a reliable model.
  • When $n$ increases further, the maximum achievable improvement of all our algorithms increases, which further demonstrates the effectiveness of our algorithms in large-scale systems.
Guideline for Algorithms Selection.
It is evident that QBC and EMCM achieved better performance than GS. However, both QBC and EMCM are model-based algorithms that require evaluating unlabeled quantum states and updating the model at each AL step. Additionally, they require training multiple models, which is computationally expensive. In contrast, GS is easy to implement and computationally inexpensive, as it simply selects the quantum state based on its geometric characteristics. Considering performance, query cost, and computational cost, we can conclude that:
  • In small-scale quantum systems or scenarios where query cost severely limits the number of queries, QBC and EMCM are generally preferable as they offer reliable performance.
  • In large-scale quantum systems, if computational cost is acceptable compared to query cost, QBC and EMCM are more suitable choices. Conversely, if computational cost is a major concern, GS is a more suitable choice.
Conclusion
In this work, we apply AL to solve a challenging problem of QPT. We propose a general framework for AL to adaptively select a subset of informative quantum states that improves the reconstruction most efficiently. In particular, we introduce a learning framework that leverages the widely-used VQCs to perform the QPT task and integrate our AL algorithms into the query step. We design and evaluate three various types of AL algorithms and provide a guideline for selecting algorithms suitable for different scenarios. Numerical results demonstrate that our algorithms achieve significantly improved reconstruction compared to the baseline that selects a subset of quantum states randomly. Moreover, these results indicate the importance of our algorithms in large-scale quantum information processing scenarios.
For future work, it will be of great interest to extend AL to other costly tasks in quantum information processing, such as shadow tomography, quantum detection tomography (QDT), quantum sensing, and quantum control (Gebhart et al. 2023).
Acknowledgements

### v2

@@ III.4 Greedy Sampling
IV Numerical Experiments
In this section, we validate our algorithms through numerical simulations on quantum processes generated by random quantum circuits. Notably, QPT has been theoretically proven to be a fundamentally difficult task [31]. As a result, most state-of-the-art studies typically perform numerical simulations on systems with maximum sizes ranging from four and seven qubits [9, 32, 33]. Therefore, we conduct simulations with up to seven qubits to evaluate the scalability of our algorithms.
IV.1 Experimental Settings
Unknown Unitary Process.— We consider the quantum process of randomly generated circuits, as shown in Fig. 2. The circuit begins and ends with Hadamard gates applied to each qubit. Each two-qubit layer is counted as a depth, including controllesd-phase (CZ) gates alternating between odd and even depths, and randomly chosen single-qubit gate ($T$, $\sqrt{X}$, or $\sqrt{Y}$). It is noted that such randomly generated quantum circuits are hard for efficient simulation on a classical computer [34, 35].
Figure 2: Structure of a randomly generated quantum circuit.
Baseline and Performance Metrics.— We consider a baseline random (RAND) that selects a subset of quantum states randomly. Additionally, we use similarity between the target process $\mathcal{U}$ and the approximate reconstruction $\mathcal{C}(\bm{\theta})$ to measure performance, defined as
$\textnormal{similarity}\left(\mathcal{U},\mathcal{C}(\bm{\theta})\right)=1-\frac{\|\mathcal{C}(\bm{\theta})-\mathcal{U}\|_{F}}{2\|\mathcal{U}\|_{F}}.$
Here, $\|\cdot\|_{F}$ denotes the Frobenius norm of a matrix. The $\text{similarity}(\mathcal{U},\mathcal{C}(\bm{\theta}))=1$ means $\|\mathcal{C}(\theta)-\mathcal{U}\|_{F}=0$, in which case $\mathcal{U}$ can be perfectly reconstructed from $\mathcal{C}(\bm{\theta})$.
@@ IV.1 Experimental Settings
Other Experimental Parameters.— We conduct simulations with different numbers of qubits to evaluate the scalability of our algorithms. For an $n$-qubit quantum process ($n$ ranges from 2 to 7), we set the VQC depth to 3, 5, 7, 8, 8, and 8 to learn the corresponding quantum processes. The number of committee members in QBC and the number of ensemble models in EMCM is empirically set to be 6. To avoid random fluctuation, each experiment is repeated 100 times and the averaged metrics are reported.
IV.2 Experimental Results
Figure 3: Similarity under different query strategies: QBC (green squares), EMCM (red triangles), GS (orange plus signs), and RAND ((blue circles). Subplots (a)–(f) correspond to 2- to 7-qubit systems, respectively.
Similarity Metric.— Fig. 3 shows the similarity comparison between our three proposed AL algorithms and the baseline RAND across varying numbers of labeled quantum states. The key observations are as follows.
@@ IV.2 Experimental Results
  • QBC and EMCM significantly outperformed RAND across all numbers of labeled quantum states. GS initially performs worse than RAND when only a few quantum states are labeled. However, as the number of labeled quantum states increases, GS eventually outperforms RAND. A possible explanation for the phenomena is as follows. At the initial steps, GS might select some outliers to maximize the distance from the previously labeled quantum states, which leads to a decrease in performance. Since the number of outliers is usually very small, their impact diminishes as more quantum states are labeled. It is worth noting that as the system size increases, reconstructing the quantum process requires a significantly larger number of labeled quantum states, which helps mitigate the impact of outliers.
Fig. 4 shows the similarity comparison between our three proposed AL algorithms and the baseline RAND. Here, the number of labeled quantum states corresponds to the minimum required for all our AL algorithms to achieve a similarity of 0.9. Similarly, we conduct simulations on quantum processes with sizes ranging from 2 to 7 qubits. It is clearly from Fig. 4 that as the system size increases, the performance gap between our AL algorithms and the baseline RAND becomes more significant. As the system size increases to 7 qubits, all AL algorithms achieve a similarity above 0.9. Meanwhile, the baseline RAND remains below 0.5.
@@
Improvement Metric.— Fig. 5 shows the improvement comparison between our three proposed AL algorithms and the baseline RAND across varying numbers of labeled quantum states. A key finding is that as the number of labeled quantum states increases, the improvement of QBC and EMCM initially increases and then decreases. This is because, with sufficient labeled quantum states, even randomly selected quantum states can lead to well-trained models. In contrast, the improvement of GS initially decreases, which is consistent with our previous analysis, as GS may select outliers in the initial steps. As more quantum states are labeled, their impact diminishes and the improvement increases then decreases, similar to QBC and EMCM.
Fig. 6 shows the improvement of our three proposed AL algorithms compared to the baseline RAND. Here, the number of labeled quantum states corresponds to the minimum required for all our AL algorithms to achieve a similarity of 0.9. Similarly, we conduct simulations on quantum processes with sizes ranging from 2 to 7 qubits. It is obvious from Fig. 6 that as the system size increases, the improvement of all AL algorithms also increases. This confirms that the improvement of our AL algorithms over the baseline RAND increases with the system size. As the system size increases to 7 qubits, the improvement of all AL algorithms exceeds 2. This indicates that, given the same number of labeled quantum states (i.e., queries to the unknown process), our AL algorithms can achieve more than twice the reconstruction similarity compared to the baseline RAND. These results further demonstrate the effectiveness of our AL algorithms, especially in large-scale quantum systems.
Figure 5: Improvement under different query strategies compared with the baseline RAND: QBC (green squares), EMCM (red triangles), and GS (orange plus signs). Subplots (a)–(f) correspond to 2- to 7-qubit systems, respectively.
[…]
Practical Guidelines for Algorithms Selection.— It is evident that QBC and EMCM achieved better performance than GS. However, both QBC and EMCM are model-based algorithms that require evaluating each unlabeled quantum states and updating the model at every AL step. Additionally, they rely on training multiple models, which is computationally expensive. In particular, QBC involves training a committee of models to quantify their disagreement in predictions, and EMCM involves training an ensemble of models to estimate the expected model change. In contrast to QBC and EMCM, GS is a model-free approach that selects quantum states based solely on their geometric characteristics. GS allows selecting a set of quantum states at the beginning of the AL process. This makes it easy to implement and computationally inexpensive since it avoids evaluating each unlabeled quantum state and updating the model at every AL step.
Considering performance, query cost, and computational cost, we can conclude that: In scenarios where query cost is the primary bottleneck (e.g., limited access to quantum hardware or high experimental costs), QBC and EMCM are generally preferable as they offer reliable performance even under a constrained query budget. Conversely, if computational cost is a major concern (e.g., limited computational resources or large-scale simulations), GS is a more suitable choice as it can achieve reasonable performance with considerably less computational cost.
V Conclusion
In this work, we apply active learning to address the challenging problem of quantum process tomography. Rather than replacing existing QPT methods, our goal is to further enhance them by optimizing the crucial query step. We propose a general AL framework that adaptively selects a subset of informative quantum states to improve the reconstruction most efficiently. We design and evaluate three different types of AL algorithms and provide guidelines for selecting suitable algorithms in various scenarios. In particular, we introduce a learning framework that leverages the widely-used VQCs to perform the QPT task and integrate our AL algorithms into the query step. We demonstrate our algorithms by reconstructing the unitary quantum processes resulting from random quantum circuits with up to seven qubits. Numerical results show that our AL algorithms achieve significantly improved reconstruction compared to the baseline method that select a subset of quantum states randomly. Moreover, the improvement increases with the size of the underlying quantum system. Most importantly, our work introduces a new perspective for advancing existing QPT methods via active learning.
For future work, since we emphasize that our AL framework is general and can in principle be applied in the query step of any QPT method, one possible direction is to generalize it to other QPT methods beyond the variational quantum circuits considered here, such as tensor network QPT [12]. Another exciting direction is to extend the concept of active learning to other costly tasks in quantum information processing, such as shadow tomography, quantum detection tomography (QDT), quantum sensing, and quantum control [36].
Acknowledgements.
This work was partially supported by the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302901), and National Natural Science Foundation of China (Grant No. 62102388).